# Optimality of quasi-Monte Carlo methods and suboptimality of the sparse-grid Gauss–Hermite rule

### v1

@@ 3 Gauss–Hermite sparse-grid quadrature
$S_{\Lambda}:=\sum_{(\ell_{1},\dots,\ell_{d})\in\Lambda}\Delta_{\ell_{1}}\otimes\dotsb\otimes\Delta_{\ell_{d}}.$ (3)
@@ 3.1 Lower bounds
Proof.
Our strategy follows [41, Section 3.1]. The heart of the matter is to construct a suitable fooling function, which we denote by $h_{\Lambda}$ below. The function is based on the earlier construction from [41, Section 3.1]. Without loss of generality, we assume that $\Lambda$ is downward-closed in the first coordinate. Then, for some $\overline{\ell}_{1}\in\mathbb{N}$ we have $(\ell_{1},1\dots,1)\in\Lambda$ for all $\ell_{1}=1,\dots,\overline{\ell}_{1}$. For example, if $\Lambda$ is the classical isotropic index set $\Lambda=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid\,|\boldsymbol{\ell}|_{1}\leq L\}$ then $\overline{\ell}_{1}=L-d+1$.
Let $n:=n_{\overline{\ell}_{1}}$ denote the number of quadrature points used in the univariate Gauss–Hermite rule $Q_{\overline{\ell}_{1}}^{\mathrm{uni}}$, and let $\xi_{1},\dots,\xi_{n}$ be the roots of the univariate Hermite polynomial $H_{n}$ of degree $n$. Following [41], define the function $p_{n}(t)\colon\mathbb{R}\to\mathbb{R}$ by
@@
$h_{\Lambda}(\boldsymbol{x}):=p_{n}(x_{1})\quad\text{for }\boldsymbol{x}=(x_{1},\dots,x_{d})\in\mathbb{R}^{d}.$ (4)
@@
$\frac{|I(h_{\Lambda})-S_{\Lambda}(h_{\Lambda})|}{\|h_{\Lambda}\|_{H_{\rho}^{\alpha}}}=\frac{\int_{\mathbb{R}}p_{n}(x)\rho(x)\mathrm{d}x}{\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}}^{2}\right)},$ (5)
@@
$\|h_{\Lambda}\|_{H_{\rho}^{\alpha}}$ $=\left(\sum_{|\mathbf{r}|_{\infty}\leq\alpha}\|D^{(r_{1},0,\dots,0)}h_{\Lambda}\|_{L_{\rho}^{2}}^{2}\right)^{\frac{1}{2}}=\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}}^{2}\right)^{\frac{1}{2}}$
@@
The right-hand side of (5) corresponds to the expression in the last displayed equation in [41, Proof of Lemma 3.1], which is further analysed in [41, Theorem 3.2]. There, it is shown
$\frac{\int_{\mathbb{R}}p_{n}(x)\rho(x)\mathrm{d}x}{\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}}^{2}\right)}\geq c_{\alpha}n^{-\alpha/2},$
@@
the statement follows. ∎
@@
$\Lambda_{L}=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid|\boldsymbol{\ell}|_{1}\leq L\},$
@@
$\Lambda_{L}^{0}=\{\boldsymbol{\ell}\in\mathbb{N}_{0}^{d}\mid|\boldsymbol{\ell}|_{1}\leq L\}$
@@
Let $S_{\Lambda_{L}}$ be the sparse-grid quadrature based on the Gauss–Hermite rule for $\Lambda_{L}=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid|\boldsymbol{\ell}|_{1}\leq L\}$. Suppose that the Gauss–Hermite rule $Q_{\ell}^{\mathrm{uni}}$ for univariate functions employs $n_{\ell}$ quadrature points for every $\ell\geq 1$, and that for some $M_{0}>0$ and $M>1$ we choose $n_{\ell}$ as
@@
Given the choice $\Lambda_{L}=\bigl\{\boldsymbol{\ell}=(\ell_{1},\dots,\ell_{d})\in\mathbb{N}^{d}:|\boldsymbol{\ell}|_{1}\leq L\bigr\}$, and in view of the well-known representation formula of the Smolyak algorithm (see e.g., [54, Lemma 15.1]) the total number of point evaluation $N$ can be written explicitly as
$N=\sum_{\boldsymbol{\ell}\in\Lambda_{L}}\prod_{k=1}^{d}n_{\ell_{k}}.$ (6)
@@
$M^{-d}\sum_{\boldsymbol{\ell}\in\Lambda_{L}}M^{|\boldsymbol{\ell}|_{1}}\leq N\leq M_{0}^{d}\sum_{\boldsymbol{\ell}\in\Lambda_{L}}M^{|\boldsymbol{\ell}|_{1}},$
and thus together with $\sum_{\boldsymbol{\ell}\in\Lambda_{L}}M^{|\boldsymbol{\ell}|_{1}}=\sum_{k=d}^{L}M^{k}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{N}^{d}\\ |\boldsymbol{\ell}|_{1}=k\end{subarray}}1$ and $\{\boldsymbol{\ell}\in\mathbb{N}^{d}:|\boldsymbol{\ell}|_{1}=k\}=\frac{(k-1)!}{(k-d)!(d-1)!}=\Theta_{d}(k^{d-1})$ we obtain
$\underline{C}_{d}M^{L}L^{d-1}\leq\underline{C}^{\prime}_{d}M^{L}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{N}^{d}\\ |\boldsymbol{\ell}|_{1}=L\end{subarray}}1\leq N\leq\overline{C}_{d}L^{d-1}\sum_{k=0}^{L}M^{k}\leq\frac{\overline{C}_{d}M}{M-1}L^{d-1}M^{L}.$ (7)
@@
which completes the proof. ∎
@@ 3.2 Application: lower bounds for sparse-grid interpolation
Following the arguments in the proofs of Theorems 3.1 and 3.2, the statement follows. ∎
@@ 3.3 Upper bound
$\Lambda_{L}:=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid\,|\boldsymbol{\ell}|_{1}\leq L\}.$
@@
To derive an upper bound, we will invoke the general theory of Smolyak algorithm [65]; see also Smolyak’s foundational work [61]. The theory therein is for general approximation problem of tensor product of continuous linear functionals defined on the tensor product of a reproducing kernel Hilbert space. Although such theory can be applied here in our setting through the norm-equivalence with the Hermite spaces, such a strategy will induce an extra constant possibly dependent on $d$. Instead, we will directly verify that an analogous result holds in $H_{\rho}^{\alpha}$. For notational convenience, define $e(S_{\Lambda}(f)):=\sup_{f\neq 0}\frac{|I(f)-S_{\Lambda}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}.$ The following result is essentially by Wasilkowski and Woźniakowski [65].
@@
Proof.
We closely follow the argument in [65, Proof of Lemma 2]. The same reasoning applies if the functional norm on $H_{\rho}^{\alpha}(\mathbb{R}^{d})$ is bounded by the product of functional norms on $H_{\rho}^{\alpha}(\mathbb{R})$. Since such a bound is established in Section 2, the result follows. ∎
Then, the final piece for deriving the numerical integration error is an error analysis in the space that gives $H_{\rho}^{\alpha}$ upon taking the tensor product. Such result is obtained in our previous paper [41].
@@
Proof.
From [41, Proposition 3.4], we know that the worst-case error in the Sobolev space for univariate functions is given by
@@
with $C_{1}:=c_{\alpha}M^{\alpha/2}$. This bound enables us to invoke Proposition 3.4, yielding
@@
To express the right-hand side above in terms of $N$, we obtain some bounds for powers of $M$. The lower bound $M^{L-d}\leq n_{L-d+1}\leq N$ can be obtained by considering the contribution for the quadrature corresponding to the index $(\ell_{1},\dots,\ell_{d})$ with $\ell_{1}=L-d+1$ and $\ell_{2}=\dotsb=\ell_{d}=1$. On the other hand, following the calculations in [65, Section 5] we get $N\leq M_{0}^{d}\frac{M}{M-1}M^{L}\binom{L-1}{d-1}$. From $\binom{L-1}{d-1}=\big(1-\frac{d-1}{L}\big)\binom{L}{d-1}\leq\binom{L}{d-1}$
@@
[…]
@@
Hence, the statement follows with $c_{\alpha,d}:=\frac{C_{2}((1+d)/\ln M)^{(d-1)(1+\alpha/2)}}{((d-1)!)^{1+\alpha/2}}\left(\frac{M_{0}^{d}M}{M-1}\right)^{\alpha/2}$. ∎
@@ 4 Quasi-Monte Carlo methods
with nodes $\boldsymbol{t}^{(j)}\in[0,1]^{d}$, $j=1,\dots,N$. Popular classes for the choice of $\boldsymbol{t}_{j}$ include lattice points [60], higher-order digital nets [14], and low discrepancy points/sequences [52, Chapter 3]. These classes are not mutually exclusive: lattice points and digital nets may also have low discrepancy; see, for instance, [12, Chapter 5] and [14, Chapter 5].
@@
$\int_{\mathbb{R}^{d}}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}$ $=\int_{\Omega}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}+\int_{\mathbb{R}^{d}\setminus\Omega}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}$ (8)
@@
$\int_{\Omega}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}$ $=\int_{[0,1]^{d}}f(\boldsymbol{\Psi}(\boldsymbol{t}))\,\rho(\boldsymbol{\Psi}(\boldsymbol{t}))\,|\det(D\boldsymbol{\Psi}(\boldsymbol{t}))|\,\mathrm{d}\boldsymbol{t}$
$\approx\frac{1}{N}\sum_{j=1}^{N}f\bigl(\boldsymbol{\Psi}(\boldsymbol{t}^{(j)})\bigr)\rho\bigl(\boldsymbol{\Psi}(\boldsymbol{t}^{(j)})\bigr)|\det(D\boldsymbol{\Psi}(\boldsymbol{t}^{(j)}))|=:Q_{N}(f).$ (9)
Here, $\Omega\subseteq\mathbb{R}^{d}$ is a suitable integration domain, possibly $\Omega=\mathbb{R}^{d}$. Often the mapping is component-wise, $\boldsymbol{\Psi}(\boldsymbol{t})=(\Psi_{1}(t_{1}),\dots,\Psi_{d}(t_{d}))$, and non-negative, in which case the Jacobian factor simplifies as
@@
In this paper, we consider two choices for $\Psi$: an affine map and a Möbius transformation. For the former, $\Omega$ is a suitable box; for the latter, $\Omega$ is the whole $\mathbb{R}^{d}$. We will show that both of them map $f\in H_{\rho}^{\alpha}$ to Sobolev spaces over the unit cube for which QMC methods are known to work well.
@@
where $\{\cdot\}$ expresses the fractional part of each component, and the integer vector $\boldsymbol{z}\in\mathbb{N}^{d}$ is called a generating vector. A good generating vector can be found efficiently using a greedy algorithm known as the component-by-component (CBC) construction, which provably yields a vector with a satisfactorily small worst-case error in various function spaces [12].
@@
A digital net is defined as follows. For a fixed prime $p$ and a positive integer $m$, we denote the $p$-adic expansion of a non-negative integer $i<p^{m}$ and a real number $x\in[0,1)$ by
$i=\iota_{0}+\iota_{1}p+\cdots\iota_{m-1}p^{m-1}\qquad\text{and}\qquad x=\frac{\xi_{1}}{p}+\frac{\xi_{2}}{p^{2}}+\cdots,$
@@
respectively. We then write $\vec{i}=(\iota_{0},\iota_{1},\ldots,\iota_{m-1})^{\top}\in\mathbb{Z}_{p}^{m}$ and $\vec{x}=(\xi_{1},\xi_{2},\ldots)^{\top}\in\mathbb{Z}_{p}^{\mathbb{N}}$. Here, $\mathbb{Z}_{p}$ is the set $\{0,\dots,p-1\}$ equipped with the usual arithmetic operations modulo $p$. If $x$ admits two different $p$-adic expansions, we choose the one in which infinitely many of the digits $\xi_{i}$ differ from p1p-1.
Now, for suitably designed generating matrices $C_{1},\ldots,C_{d}\in\mathbb{Z}_{p}^{\mathbb{N}\times m}$, a digital net consisting of $N=p^{m}$ points is given by
$\vec{t}_{j,i}=C_{i}\,\overrightarrow{(j-1)}\;\;\text{ for }i=1,\ldots,d\text{ and }j=1,\ldots,N.$
@@
Note that if every column of the generating matrices has only finitely many non-zero entries, this matrix-vector product is well-defined. The classical (first-order) digital nets constructed by Sobol’ [62], Niederreiter [50], and Niederreiter–Xing [51] are included as special cases of this construction. In [10], Dick introduced the notion of higher-order digital nets, which also form a special class within this framework. An order-$\alpha$ digital net with $\alpha\in\mathbb{N}$, $\alpha\geq 2$, is constructed using generating matrices $C_{1},\ldots,C_{d}$, which are themselves derived from the generating matrices $D_{1},\ldots,D_{\alpha d}$ for first-order digital nets in dimension $\alpha d$. We refer the reader to [28] for a review of higher-order nets.
@@
It is worth noting that QMC methods presented in this section always have positive weights, i.e., when seeing the quadrature (9) as $Q_{N}(f)=\sum_{j=1}^{N}w_{j}f(\boldsymbol{x}_{j})$, weights $w_{j}$’s are all positive. As such, nonnegative integrands yield nonnegative approximated integral values. This nonnegativity-preserving property is important in situations where the quantity of interest must be nonnegative, for example probabilities. Positive quadrature weights are also useful when discretizing optimization problems involving expectations [46] with convex objective functions, since they preserve convexity. By contrast, other quadrature methods may involve negative weights and thus lose these desirable features.

### v2

@@ 3 Gauss–Hermite sparse-grid quadrature
(3) $S_{\Lambda}:=\sum_{(\ell_{1},\dots,\ell_{d})\in\Lambda}\Delta_{\ell_{1}}\otimes\dotsb\otimes\Delta_{\ell_{d}}.$
@@ 3.1 Lower bounds
Proof 3.2.
Our strategy follows [undefap, Section 3.1]. The heart of the matter is to construct a suitable fooling function, which we denote by $h_{\Lambda}$ below. Without loss of generality, we assume that $\Lambda$ is downward-closed in the first coordinate. Then, for some $\overline{\ell}_{1}\in\mathbb{N}$ we have $(\ell_{1},1\dots,1)\in\Lambda$ for all $\ell_{1}=1,\dots,\overline{\ell}_{1}$. For example, if $\Lambda$ is the classical isotropic index set $\Lambda=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid\,|{\boldsymbol{\ell}|}\leq L\}$ then $\overline{\ell}_{1}=L-d+1$.
Let $n:=n_{\overline{\ell}_{1}}$ denote the number of quadrature points used in the univariate Gauss–Hermite rule $Q_{\overline{\ell}_{1}}^{\mathrm{uni}}$, and let $\xi_{1},\dots,\xi_{n}$ be the roots of the univariate Hermite polynomial $H_{n}$ of degree $n$. Following [undefap], define the function $p_{n}(t)\colon\mathbb{R}\to\mathbb{R}$ by
@@
(4) $h_{\Lambda}(\boldsymbol{x}):=p_{n}(x_{1})\quad\text{for }\boldsymbol{x}=(x_{1},\dots,x_{d})\in\mathbb{R}^{d}.$
@@
(5) $\frac{|I(h_{\Lambda})-S_{\Lambda}(h_{\Lambda})|}{\|h_{\Lambda}\|_{H_{\rho}^{\alpha}}}=\frac{\int_{\mathbb{R}}p_{n}(x)\rho(x)\mathrm{d}x}{\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}}^{2}\right)^{{1/2}}},$
@@
$\|h_{\Lambda}\|_{H_{\rho}^{\alpha}}$ $=\left(\sum_{|\mathbf{r}|_{\infty}\leq\alpha}\|D^{(r_{1},0,\dots,0)}h_{\Lambda}\|_{L_{\rho}^{2}}^{2}\right)^{\frac{1}{2}}=\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}}^{2}\right)^{\frac{1}{2}}{.}$
@@
The right-hand side of (5) corresponds to the expression in the last displayed equation in [undefap, Proof of Lemma 3.1], which is further analysed in [undefap, Theorem 3.2]. There, it is shown
$\frac{\int_{\mathbb{R}}p_{n}(x)\rho(x)\mathrm{d}x}{\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}}^{2}\right)^{{1/2}}}\geq c_{\alpha}n^{-\alpha/2},$
@@
where #(Q)\#(Q) denotes the total number of quadrature nodes used in a quadrature rule $Q$, the statement follows.
@@
$\Lambda_{L}=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid|{\boldsymbol{\ell}|}\leq L\},\quad{\text{where }\ |\boldsymbol{\ell}|=\ell_{1}+\dots+\ell_{d},}$
@@
$\Lambda_{L}^{0}=\{\boldsymbol{\ell}\in\mathbb{N}_{0}^{d}\mid|{\boldsymbol{\ell}|}\leq L\}$
@@
Let $S_{\Lambda_{L}}$ be the sparse-grid quadrature based on the Gauss–Hermite rule for $\Lambda_{L}=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid|{\boldsymbol{\ell}|}\leq L\}$. Suppose that the Gauss–Hermite rule $Q_{\ell}^{\mathrm{uni}}$ for univariate functions employs $n_{\ell}$ quadrature points for every $\ell\geq 1$, and that for some $M_{0}>0$ and $M>1$ we choose $n_{\ell}$ as
@@
Given the choice $\Lambda_{L}=\bigl\{\boldsymbol{\ell}=(\ell_{1},\dots,\ell_{d})\in\mathbb{N}^{d}:|{\boldsymbol{\ell}|}\leq L\bigr\}$, and in view of the well-known representation formula of the Smolyak algorithm (see e.g., [undefaad, Lemma 15.1]) the total number of point evaluation $N$ can be written explicitly as
(6) $N=\sum_{\boldsymbol{\ell}\in\Lambda_{L}}\prod_{k=1}^{d}n_{\ell_{k}}.$
@@
$M^{-d}\sum_{\boldsymbol{\ell}\in\Lambda_{L}}M^{|{\boldsymbol{\ell}|}}\leq N\leq M_{0}^{d}\sum_{\boldsymbol{\ell}\in\Lambda_{L}}M^{|{\boldsymbol{\ell}|}},$
and thus together with $\sum_{\boldsymbol{\ell}\in\Lambda_{L}}M^{|{\boldsymbol{\ell}|}}=\sum_{k=d}^{L}M^{k}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{N}^{d}\\ |{\boldsymbol{\ell}|}=k\end{subarray}}1$ and $\{\boldsymbol{\ell}\in\mathbb{N}^{d}:|{\boldsymbol{\ell}|}=k\}=\frac{(k-1)!}{(k-d)!(d-1)!}=\Theta_{d}(k^{d-1})$ we obtain
(7) $\underline{C}_{d}M^{L}L^{d-1}\leq\underline{C}^{\prime}_{d}M^{L}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{N}^{d}\\ |{\boldsymbol{\ell}|}=L\end{subarray}}1\leq N\leq\overline{C}_{d}L^{d-1}\sum_{k=0}^{L}M^{k}\leq\frac{\overline{C}_{d}M}{M-1}L^{d-1}M^{L}.$
@@
which completes the proof.
@@ 3.2 Application: lower bounds for sparse-grid interpolation
Following the arguments in the proofs of Theorems 3.1 and 3.3, the statement follows.
@@ 3.3 Upper bound
$\Lambda_{L}:=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid\,|{\boldsymbol{\ell}|}\leq L\}.$
@@
To derive an upper bound, we will invoke the general theory of Smolyak algorithm [undefaao]; see also Smolyak’s foundational work [undefaak]. The theory therein is for general approximation problem of tensor product of continuous linear functionals defined on the tensor product of a reproducing kernel Hilbert space. Although such theory can be applied here in our setting through the norm-equivalence with the Hermite spaces, such a strategy will induce an extra constant possibly dependent on $d$. Instead, we will directly verify that an analogous result holds in $H_{\rho}^{\alpha}$. For notational convenience, define $e(S_{\Lambda}(f)):=\sup_{f\neq 0}\frac{|I(f)-S_{\Lambda}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}.$ The following result is essentially by Wasilkowski and Woźniakowski [undefaao].
@@
with $C_{2}:=C_{1}\Bigl(\max\bigl\{M^{\alpha/2}\,,\,{C_{1}(1+M^{\alpha/2})}\bigr\}\Bigr)^{d-1}$.
Proof 3.8.
We closely follow the argument in [undefaao, Proof of Lemma 2]. The same reasoning applies if bounded linear functionals on $H_{\rho}^{\alpha}(\mathbb{R}^{s})$ with $\rho(\boldsymbol{x})=\frac{\mathrm{e}^{-|\boldsymbol{x}|_{2}^{2}/2}}{(2\pi)^{s/2}}$ of the form $T_{1}\otimes\dotsb\otimes T_{s}$, where each $T_{j}\colon H_{\rho}^{\alpha}(\mathbb{R})\to\mathbb{R}$ is a bounded linear funcitonal on $H_{\rho}^{\alpha}(\mathbb{R})$ with $\rho(x)=\frac{\mathrm{e}^{-|x|^{2}/2}}{(2\pi)^{1/2}}$, satisfies $\|T_{1}\otimes\dotsb\otimes T_{s}\|_{H_{\rho}^{\alpha}(\mathbb{R}^{s})\to\mathbb{R}}=\prod_{j=1}^{s}\|T_{j}\|_{H_{\rho}^{\alpha}(\mathbb{R})\to\mathbb{R}}$ for $s=1,\dots,d$. Since such an equality holds as per Section 2.2, the result follows.
Then, the final piece for deriving the numerical integration error is an error analysis in the space that gives $H_{\rho}^{\alpha}$ upon taking the tensor product. Such result is obtained in our previous paper [undefap].
@@
Proof 3.10.
From [undefap, Proposition 3.4], we know that the worst-case error in the Sobolev space for univariate functions is given by
@@
with a constant $c_{\alpha}$ depending on $\alpha$, and $C_{1}:=c_{\alpha}M^{\alpha/2}$. This bound enables us to invoke Proposition 3.7, yielding
@@
To express the right-hand side above in terms of $N$, we obtain some bounds for powers of $M$. The lower bound $M^{L-d}\leq n_{L-d+1}\leq N$ can be obtained by considering the contribution for the quadrature corresponding to the index $(\ell_{1},\dots,\ell_{d})$ with $\ell_{1}=L-d+1$ and $\ell_{2}=\dotsb=\ell_{d}=1$. On the other hand, following the calculations in [undefaao, Section 5] we get $N\leq M_{0}^{d}\frac{M}{M-1}M^{L}\binom{L-1}{d-1}$. From $\binom{L-1}{d-1}=\big(1-\frac{d-1}{L}\big)\binom{L}{d-1}\leq\binom{L}{d-1}$
@@
To bound the factor of $L$ in terms of $\ln N$, note that $2\leq L-d+1$ implies $M\leq n_{2}\leq N$; combined with $M^{L-d}\leq N$, this gives $M^{L}\leq M^{d}N\leq N^{1+d}$ and thus $L\leq{(}\ln N{)}(1+d)/\ln M$. Therefore, we obtain
@@
Hence, the statement follows with $c_{\alpha,d}:=\frac{C_{2}((1+d)/\ln M)^{(d-1)(1+\alpha/2)}}{((d-1)!)^{1+\alpha/2}}\left(\frac{M_{0}^{d}M}{M-1}\right)^{\alpha/2}$.
@@ 4 Quasi-Monte Carlo methods
with nodes $\boldsymbol{t}^{(j)}\in[0,1]^{d}$, $j=1,\dots,N$. Popular classes for the choice of $\boldsymbol{t}_{j}$ include lattice points [undefaab, undefaaj, undefl], (higher-order) digital nets [undefaab, undefn], and low discrepancy points/sequences (such as Halton and Kronecker sequences) [undefaab, Chapter 3]. These classes are not mutually exclusive: lattice points and digital nets may also have low discrepancy; see, for instance, [undefl, Chapter 5] and [undefn, Chapter 5].
@@
(8) $\int_{\mathbb{R}^{d}}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}$ $=\int_{\Omega}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}+\int_{\mathbb{R}^{d}\setminus\Omega}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}$
@@
$\int_{\Omega}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}$ $=\int_{{(}0,1{)}^{d}}f(\boldsymbol{\Psi}(\boldsymbol{t}))\,\rho(\boldsymbol{\Psi}(\boldsymbol{t}))\,|\det(D\boldsymbol{\Psi}(\boldsymbol{t}))|\,\mathrm{d}\boldsymbol{t}$
(9) $\approx\frac{1}{N}\sum_{j=1}^{N}f\bigl(\boldsymbol{\Psi}(\boldsymbol{t}^{(j)})\bigr)\rho\bigl(\boldsymbol{\Psi}(\boldsymbol{t}^{(j)})\bigr)|\det(D\boldsymbol{\Psi}(\boldsymbol{t}^{(j)}))|=:Q_{N}(f).$
Here, $\Omega\subseteq\mathbb{R}^{d}$ is a suitable integration domain, possibly $\Omega=\mathbb{R}^{d}$, and $\boldsymbol{\Psi}:(0,1)^{d}\to\Omega$ is a suitable mapping. Often the mapping is component-wise, $\boldsymbol{\Psi}(\boldsymbol{t})\!\!\;=\!\!\;(\Psi_{1}(t_{1}),\dots,\Psi_{d}(t_{d}))$, and non-negative, in which case the Jacobian factor simplifies as
@@
In this paper, we consider two choices for $\Psi$: an affine map and a Möbius transformation, namely the cotangent transform. For the former, $\Omega$ is a suitable box; for the latter, $\Omega$ is the whole $\mathbb{R}^{d}$. We will show that both of them map $f\in H_{\rho}^{\alpha}$ to Sobolev spaces over the unit cube for which QMC methods are known to work well.
@@
where $\{\cdot\}$ expresses the fractional part of each component, and the integer vector $\boldsymbol{z}\in\mathbb{N}^{d}$ is called a generating vector. A good generating vector can be found efficiently using a greedy algorithm known as the component-by-component (CBC) construction, which provably yields a vector with a satisfactorily small worst-case error in various function spaces [undefl].
@@
A digital net is defined as follows. For a fixed prime $p$ and a positive integer $m$, we denote the $p$-adic expansion of a non-negative integer ${j}<p^{m}$ and a real number $x\in[0,1)$ by
${j=\iota_{0}+\iota_{1}p+\cdots+\iota_{m-1}p^{m-1}}\qquad\text{and}\qquad x=\frac{\xi_{1}}{p}+\frac{\xi_{2}}{p^{2}}+\cdots,$
@@
respectively. Here, the digits $\iota_{k}$ and $\xi_{k}$ are in $\mathbb{Z}_{p}$, the set $\{0,1,\dots,p-1\}$ equipped with the usual arithmetic operations modulo $p$. We then write $\vec{j}=(\iota_{0},\iota_{1},\ldots,\iota_{m-1})^{\top}\in\mathbb{Z}_{p}^{m}$ and $\vec{x}=(\xi_{1},\xi_{2},\ldots)^{\top}\in\mathbb{Z}_{p}^{\mathbb{N}}$. If $x$ admits two different $p$-adic expansions, we choose the one in which infinitely many of the digits $\xi_{k}$ differ from p1p-1.
Now, for suitably designed generating matrices $C_{1},\ldots,C_{d}\in\mathbb{Z}_{p}^{\mathbb{N}\times m}$, consider the corresponding digit vectors $\vec{t}_{j,i}\in\mathbb{Z}^{\mathbb{N}}_{p}$ obtained by the matrix-vector product modulo $p$:
$\vec{t}_{j,i}=C_{i}\,\vec{j}\;\;\text{ for }i=1,\ldots,d\text{ and }j=0,1,\ldots,N-1.$
@@
Then, the digital net consisting of $N=p^{m}$ points corresponding to $C_{1},\dots,C_{d}$ is defined as the set $\{\boldsymbol{t}^{(0)},\ldots,\boldsymbol{t}^{(N-1)}\}$, where each component $t_{j,i}$ of the point $\boldsymbol{t}^{(j)}=(t_{j,1},\ldots,t_{j,d})\in[0,1]^{d}$ is determined by the corresponding digit vector $\vec{t}_{j,i}$. Note that if every column of the generating matrices has only finitely many non-zero entries, this matrix-vector product is well-defined. The classical (first-order) digital nets constructed by Sobol’ [undefaal], Niederreiter [undefaz], and Niederreiter–Xing [undefaaa] are included as special cases of this construction. In [undefj], Dick introduced the notion of higher-order digital nets, which also form a special class within this framework. An order-$\alpha$ digital net with $\alpha\in\mathbb{N}$, $\alpha\geq 2$, is constructed using generating matrices $C_{1},\ldots,C_{d}$, which are themselves derived from the generating matrices $D_{1},\ldots,D_{\alpha d}$ for first-order digital nets in dimension $\alpha d$. We refer the reader to [undefab] for a review of higher-order nets.
@@
It is worth noting that QMC methods presented in this section always have positive weights, i.e., when seeing the quadrature (9) as $Q_{N}(f)=\sum_{j=1}^{N}w_{j}f(\boldsymbol{x}_{j})$, weights $w_{j}$’s are all positive. As such, nonnegative integrands yield nonnegative approximated integral values. This nonnegativity-preserving property is important in situations where the quantity of interest must be nonnegative, for example probabilities. Positive quadrature weights are also useful when discretizing optimization problems involving expectations [undefav] with convex objective functions, since they preserve convexity. By contrast, other quadrature methods may involve negative weights and thus lose these desirable features.
Before going into details, we first summarize our main results of this section in the following theorem.
Theorem 4.1.
The quasi-Monte Carlo rule $Q_{N}$ (9) achieves the error bound
where the exponent $\gamma$ depends on the QMC construction:
[…]
The bound for higher-order digital nets with an affine map is due to Dick et al. [undefk]. In the following subsections, we present detailed algorithms and establish the remaining upper bounds. The optimality statements then follow from the lower bound in Theorem 2.1.